# VecFontSDF: Learning to Reconstruct and Synthesize High-quality Vector Fonts via Signed Distance Functions

### v3

@@ 1 Introduction
Traditional vector font designing process relies heavily on the expertise and effort from professional designers, setting a high barrier for common users. With the rapid development of deep generative models in the last few years, a large amount of effective and powerful methods [1, 31, 6] have been proposed to synthesize visually-pleasing glyph images. In the meantime, how to automatically reconstruct and generate high-quality vector fonts is still considered as a challenging task in the communities of Computer Vision and Computer Graphics. Recently, several methods based on sequential generative models [carlier2020depdfvg, 10, 19, 29, 32] have been reported that treat a vector glyph as a draw-commands sequence and use Recurrent Neural Networks (RNNs) or Transformer [30] to encode and decode the sequence. However, this explicit representation of vector graphics is extremely difficult for the learning and comprehension of deep neural networks, mainly due to the long-range dependence and the ambiguity in how to draw the outlines of glyphs. More recently, DeepVecFont [32] was proposed to use dual-modality learning to alleviate the problem, showing state-of-the-art performance on this task. Its key idea is to use a CNN encoder and an RNN encoder to extract features from both image and sequence modalities. Despite using richer information of dual-modality data, it still needs repetitively random samplings of synthesis results to find the optimal one and then uses Diffvg [17] to refine the vector glyphs under the guidance of generated images in the inference stage.
Another possible solution to model the vector graphic is to use implicit functions which have been widely used to represent 3D shapes in recent years. For instance, DepdfDF [24] adopts a neural network to predict the values of signed distance functions for surfaces, but it fails to convert those SDF values to the explicit representation. BSP-Net [3] uses hyperplanes to split the space in 2D shapes and 3D meshes, but it generates unsmooth and disordered results when handling glyphs consisting of numerous curves.
@@
The work most related to ours is [18], which also aims to provide an implicit shape representation for glyphs, but possesses a serious limitation mentioned below. For convenience, we name the representation proposed in [18] IGSR, standing for “Implicit Glyph Shape Representation”. The quadratic curves used in IGSR are not strictly parabolic curves, which cannot be translated to quadratic Bézier curves. As a consequence, their model is capable of synthesizing high-resolution glyph images but not vector glyphs. Furthermore, it only uses raster images for supervision which inevitably leads to inaccurate reconstruction results. On the contrary, our proposed VecFontSDF learns to reconstruct and synthesize high-quality vector fonts that consist of quadratic Bézier curves by training on the corresponding vector data in an end-to-end manner. Major contributions of our paper are threefold:
@@ 2.1 Vector Font Generation
In early years, researchers tried to generate glyph images and utilized traditional vectorization methods [16, 23] to obtain vector fonts. With the development of sequential models such as long short-term memory RNNs [12] and Transformer [30], a number of methods have been developed that treat vector graphics as drawing-command sequences for modeling. Lopes et al. [19] proposed SVG-VAE that provides a scale-invariant representation for imagery and then uses an RNN decoder to generate vector glyphs. Carlier et al.[carlier2020depdfvg] proposed a novel hierarchical generative network called DepdfVG, which effectively disentangles high-level shapes from the low-level commands that encode the shape itself and directly predicts a set of shapes in a non-autoregressive fashion. Reddy et al.[25] proposed Im2Vec that can generate complex vector graphics with varying topologies, and only requires indirect supervision from raster images. Mo et al.[21] introduced a general framework to produce line drawings from a wide variety of images by using a dynamic window. At the same time, Wang et al. [32] adopted dual-modality learning to synthesize visually pleasing vector glyphs. However, all the methods mentioned above are based on the explicit representation of vector fonts, which requires the neural networks to learn long-range dependence and generate the drawing commands step by step. On the contrary, our method regards the vector glyph as a whole and is trained by the deterministic values of signed distance functions (SDFs).
@@ 2.2 Implicit Shape Representation
In the last two decades, lots of distance functions-based algorithms have been developed for implicit shape representation. As a pioneering work, Gibson et al. [7] proposed an adaptively sampled distance field using octree for both 2D glyph representation and 3D rendering. Green [8] used distance functions for glyph rasterization. With the rapid development of machine learning techniques, signed distance functions and other implicit functions-based methods have been widely used in 2D and 3D reconstruction tasks. Park et al. [24] proposed DepdfDF which uses an auto-decoder to predict the values of signed distance functions for 3D reconstruction. Chen et al. [4] presented an implicit function using fully-connected layers that output signed distances with the input of image features and sampling points. However, the implicit representations they used can not be directly converted to traditional explicit descriptions.
@@ 3 Method Description
In order to directly reconstruct and synthesize vector fonts from input glyph images via neural networks, we need a learnable shape representation that precisely preserves geometry information. Previous methods that fully rely on sequential generative models such as SVG-VAE [19] and DeepVecFont [32] fail to perfectly handle this task mainly due to the ambiguity of control point selection. To address this problem, we propose a vector graphic reconstructor based on an interpretable distance function with differentiable rendering. Unlike most SDF-based methods such as DepdfDF [28] which directly predicts the values of signed distance functions via neural networks, our model outputs the learnable parameters of curves that can be applied to handle complex glyph outlines containing concave curves.

### v4

@@ 1 Introduction
Traditional vector font designing process relies heavily on the expertise and effort from professional designers, setting a high barrier for common users. With the rapid development of deep generative models in the last few years, a large amount of effective and powerful methods [1, 32, 7] have been proposed to synthesize visually-pleasing glyph images. In the meantime, how to automatically reconstruct and generate high-quality vector fonts is still considered as a challenging task in the communities of Computer Vision and Computer Graphics. Recently, several methods based on sequential generative models [3, 11, 20, 30, 33] have been reported that treat a vector glyph as a draw-commands sequence and use Recurrent Neural Networks (RNNs) or Transformer [31] to encode and decode the sequence. However, this explicit representation of vector graphics is extremely difficult for the learning and comprehension of deep neural networks, mainly due to the long-range dependence and the ambiguity in how to draw the outlines of glyphs. More recently, DeepVecFont [33] was proposed to use dual-modality learning to alleviate the problem, showing state-of-the-art performance on this task. Its key idea is to use a CNN encoder and an RNN encoder to extract features from both image and sequence modalities. Despite using richer information of dual-modality data, it still needs repetitively random samplings of synthesis results to find the optimal one and then uses Diffvg [18] to refine the vector glyphs under the guidance of generated images in the inference stage.
Another possible solution to model the vector graphic is to use implicit functions which have been widely used to represent 3D shapes in recent years. For instance, DeepSDF [25] adopts a neural network to predict the values of signed distance functions for surfaces, but it fails to convert those SDF values to the explicit representation. BSP-Net [4] uses hyperplanes to split the space in 2D shapes and 3D meshes, but it generates unsmooth and disordered results when handling glyphs consisting of numerous curves.
@@
The work most related to ours is [19], which also aims to provide an implicit shape representation for glyphs, but possesses a serious limitation mentioned below. For convenience, we name the representation proposed in [19] “IGSR”, standing for “Implicit Glyph Shape Representation”. The quadratic curves used in IGSR are not strictly parabolic curves, which cannot be translated to quadratic Bézier curves. As a consequence, their model is capable of synthesizing high-resolution glyph images but not vector glyphs. Furthermore, it only uses raster images for supervision which inevitably leads to inaccurate reconstruction results. On the contrary, our proposed VecFontSDF learns to reconstruct and synthesize high-quality vector fonts that consist of quadratic Bézier curves by training on the corresponding vector data in an end-to-end manner. Major contributions of our paper are threefold:
@@ 2.1 Vector Font Generation
In early years, researchers tried to generate glyph images and utilized traditional vectorization methods [17, 24] to obtain vector fonts. With the development of sequential models such as long short-term memory RNNs [13] and Transformer [31], a number of methods have been developed that treat vector graphics as drawing-command sequences for modeling. Lopes et al. [20] proposed SVG-VAE that provides a scale-invariant representation for imagery and then uses an RNN decoder to generate vector glyphs. Carlier et al.[3] proposed a novel hierarchical generative network called DeepSVG, which effectively disentangles high-level shapes from the low-level commands that encode the shape itself and directly predicts a set of shapes in a non-autoregressive fashion. Reddy et al.[26] proposed Im2Vec that can generate complex vector graphics with varying topologies, and only requires indirect supervision from raster images. Mo et al.[22] introduced a general framework to produce line drawings from a wide variety of images by using a dynamic window. At the same time, Wang et al. [33] adopted dual-modality learning to synthesize visually pleasing vector glyphs. However, all the methods mentioned above are based on the explicit representation of vector fonts, which requires the neural networks to learn long-range dependence and generate the drawing commands step by step. On the contrary, our method regards the vector glyph as a whole and is trained by the deterministic values of signed distance functions (SDFs).
@@ 2.2 Implicit Shape Representation
In the last two decades, lots of distance functions-based algorithms have been developed for implicit shape representation. As a pioneering work, Gibson et al. [8] proposed an adaptively sampled distance field using octree for both 2D glyph representation and 3D rendering. Green [9] used distance functions for glyph rasterization. With the rapid development of machine learning techniques, signed distance functions and other implicit functions-based methods have been widely used in 2D and 3D reconstruction tasks. Park et al. [25] proposed DeepSDF which uses an auto-decoder to predict the values of signed distance functions for 3D reconstruction. Chen et al. [5] presented an implicit function using fully-connected layers that output signed distances with the input of image features and sampling points. However, the implicit representations they used can not be directly converted to traditional explicit descriptions.
@@ 3 Method Description
In order to directly reconstruct and synthesize vector fonts from input glyph images via neural networks, we need a learnable shape representation that precisely preserves geometry information. Previous methods that fully rely on sequential generative models such as SVG-VAE [20] and DeepVecFont [33] fail to perfectly handle this task mainly due to the ambiguity of control point selection. To address this problem, we propose a vector graphic reconstructor based on an interpretable distance function with differentiable rendering. Unlike most SDF-based methods such as DeepSDF [29] which directly predicts the values of signed distance functions via neural networks, our model outputs the learnable parameters of curves that can be applied to handle complex glyph outlines containing concave curves.